what is the ubia of the following assets a vehicle purchased for $30,000 and used 80% for business and apartment building purchased for $100,000 including $20,000 for the land used 100% for business and future purchase at $10,000 and used 100% for business

the correct answer is 114,000 a
Answer #1

Solution:

For purposes of determining your UBIA for all qualified property, the unadjusted basis immediately after acquisition means the basis on the placed-in-service date. Qualified property includes tangible property subject to depreciation under section 167 that is held, and used in the production of QBI, by the trade or business (or aggregated trades or businesses) during and at the close of the tax year, for which the depreciable period hasn’t ended before the close of the tax year. The depreciable period ends on the later of 10 years after the property is first placed in service by you or the last day of the last full year in the applicable recovery period under section 168(c). Additional first-year depreciation under section 168 doesn’t affect the applicable recovery period.

1. Vehicle (80 % used for business purpose) = ($ 30000* 80 %) = $ 24000

2 Building (excluding land, 100 % used for business) = ($ 100000-20000)*100 % = $ 80000

3. Future (100 % used for Business ) = ($ 10000*100%)= $ 10000

_________

$ 114000